# A systematic assessment of Large Language Models for constructing two-level fractional factorial designs

### v1

@@ Abstract
Two-level fractional factorial designs permit the study multiple factors using a limited number of runs. Traditionally, these designs are obtained from catalogs available in standard textbooks or statistical software. However, modern Large Language Models (LLMs) can now produce two-level fractional factorial designs, but the quality of these designs has not been previously assessed. In this paper, we perform a systematic evaluation of two popular classes of LLMs, namely GPT and Gemini models, to construct two-level fractional factorial designs with 8, 16, and 32 runs, and 4 to 26 factors. To this end, we use prompting techniques to develop a high-quality set of design construction tasks for the LLMs. We compare the designs obtained by the LLMs with the best-known designs in terms of resolution and minimum aberration criteria. We show that the LLMs can effectively construct optimal 8-, 16-, and 32-run designs with up to eight factors.
@@ 1 Introduction
Two-level fractional factorial designs (wu2011experiments) are commonly used experimental plans that can study many factors using an economical number of runs. In these designs, all factors are set at two levels, allowing the study of their main effects and interactions. We denote a two-level fractional factorial design as $2^{m-p}$, where $m$ is the number of factors and $p$ is the number of generated factors. The generated factors are obtained from the interactions of b=mpb=m-p basic factors, whose $2^{b}$ level combinations are all in the design (wu2011experiments). Design $2^{m-p}$ is thus a $2^{p}$ fraction of the $2^{m}$ full factorial design, with a run size that is a power of two.
@@
The DoE literature has complete catalogs of two-level fractional factorial designs. For instance, chen1993catalogue provide a catalog of designs with 8, 16, 32, and 64 runs, and up to 32 factors. ryan2010minimum provide a catalog of 128-run designs with up to 64 factors. The best designs in these catalogs are available in standard DoE textbooks such as wu2011experiments and montgomery2017design, or statistical software such as JMP, Minitab, and the FrF2 package (gromping2014r) in R.
@@
In recent years, generative artificial intelligence (GenAI) chatbots, such as ChatGPT (chatgpt2025) and Gemini (GoogleGemini), have emerged as disruptive technology that has transformed the way we learn, write, access information, make decisions, and even conduct research. For example, GenAI chatbots have been used to solve optimization problems (bertsimas2024robust), prove mathematical theorems (frieder2023mathematical), teach statistics (ellis2023new), generate code for data analyses (song2025performance), and build search engine software for statistical quality control (megahed2024introducing). The success of GenAI chatbots is due to their use of Large Language Models (LLMs; alammar2024hands) and the high computational power available these days (top500). Technically, an LLM is a deep neural network (goodfellow2016deep) with billions of parameters that are estimated through a training process that uses an enormous amount of text data. These data are typically extracted from the Internet and may include books, blogs, research and news articles, social media, among others. The goal of the training process is to predict the next word in a sentence. To this end, the LLM uses a neural network architecture called transformer (vaswani2017attention), which allows it to learn the context and relationships between words in the text during its training.
@@
In particular, ChatGPT and Gemini use the Generative Pre-trained Transformer (GPT; chatgpt2025) model and the Gemini model (comanici2025gemini), respectively, which are sophisticated, pre-trained transformer-based LLMs. To perform a specific task, their main input (as well as of other LLMs) is a user-specified text called prompt (schulhoff2024prompt), which describes the question, instruction, or guidelines for it. At the time of writing this article, the versions of the LLMs used by default in ChatGPT’s and Gemini’s web interfaces are GPT-5.1 and Gemini 2.5 Flash, respectively. Remarkably, these LLMs can reason across audio, vision, and text in real time. Moreover, they can be accessed using application programming interfaces (APIs) for experimentation.
@@
ChatGPT’s output 1 shows a $2^{4-1}$ design that matches the 4-factor 8-run design in Table 8.14 of montgomery2017design. The four factors are labeled $A$, $B$, $C$, and $D$, where the first three are basic factors and the last is the generated factor obtained as $D=ABC$. That is, the column for $D$ is the element-wise product of the columns that accommodate $A$, $B$, and $C$, when their levels are coded as 1-1 and $1$. The design matrix is shown in the Design Table section. Its first column shows labels for the runs and the other columns show the factors, whose levels are labeled using the symbols ‘-’ and ‘++’. This $2^{4-1}$ design can estimate the main effects and two-factor interactions of the four factors simultaneously. In fact, it is the best design in terms of the resolution and minimum aberration criteria introduced in Section 2. The Appendix shows a similar interaction between us and Gemini with the Gemini 2.5 Flash model, in which the same $2^{4-1}$ design was produced.
Our interactions with ChatGPT and Gemini show that their default LLMs have the potential to construct two-level fractional factorial designs. Therefore, practitioners may feel tempted to construct these designs from scratch using the chatbots, bypassing statistical software and textbooks. To explore the capabilities and limitations of this approach, we conducted a systematic study to evaluate the performance of the GPT-5.1 and Gemini 2.5 Flash models on 36 tasks concerning the construction of two-level fractional factorial designs. The tasks involve designs with 8, 16, and 32 runs, and 4 to 26 factors. Each task has a prompt to instruct the LLM to construct a design. To this end, we develop a prompt template using prompting techniques such as role, context, chain of thought, and output format (kojima2022large; white2023prompt; schulhoff2024prompt). Our prompt template belongs to the class of zero-shot prompts, introduced in Section 3, and is intended to test the general knowledge of LLMs on constructing two-level fractional factorial designs without examples.
We evaluate the computational performance of the LLMs using 10 replicates of each task. We show that the LLMs have the potential to construct optimal designs with up to 15 factors in terms of the criteria in Section 2. However, the Gemini 2.5 Flash model is better than the GPT-5.1 model because it has a high consistency in obtaining optimal 8-run designs, almost all optimal 16-run designs, and the optimal 32-run 6-factor design. For designs of other sizes, the performance of both LLMs deteriorates because they cannot consistently construct the optimal design or cannot construct a design at all.
@@ 2 Criteria for two-level fractional factorial designs
We review resolution, minimum aberration, and minimum moment aberration (wu2011experiments; xu2003minimum) to evaluate two-level fractional factorial designs. We illustrate these criteria using a $2^{7-3}$ design. Its basic factors are $A$, $B$, $C$, and $D$, and the generated factors are obtained as $E=ABC$, $F=ABD$, and $G=ACD$. The defining relation of the design is
@@ 2.1 Resolution and minimum aberration
The resolution is the length of the shortest word in the defining relation (wu2011experiments). It allows us to know the properties of a $2^{m-p}$ design for studying main effects and two-factor interactions, which are the most important according to the effect hierarchy principle (wu2011experiments, ch 4). For instance, resolution-3 designs provide main effects that are not aliased with each other, but at least one main effect is aliased with one or more two-factor interactions. In contrast, resolution-4 designs provide main effects that are not aliased with each other nor with two-factor interactions. However, at least one pair of two-factor interactions is aliased in these designs. Resolution-5 designs are optimal for studying main effects and two-factor interactions because they avoid aliasing among these effects. So, when comparing designs in terms of their resolutions, the design with the higher resolution is preferred. From Equation (1), we have that the $2^{7-3}$ design has a resolution of four.
The minimum aberration criterion allows us to differentiate between designs with the same resolution. It is based on the word length pattern (WLP; wu2011experiments), which is a vector that collects the number of words of length 1 to $m$ in the defining relation. Specifically, this vector is $(A_{1},A_{2},\ldots,A_{m})$ where $A_{i}$ is the number of words of length $i$ in the defining relation. Note that the resolution of a design is the smallest value of $i$ for which $A_{i}>0$. From Equation (1), we have that the WLP of the $2^{7-3}$ design is $(0,0,0,7,0,0,0)$. So, the defining relation of this design has seven words of length four.
The minimum aberration criterion is to sequentially minimize the WLP from left to right. In this way, the criterion favors designs with less aliasing involving low-order effects of the factors than high-order ones. chen1993catalogue show that the $2^{7-3}$ design in Table 1 has minimum aberration among all two-level designs of the same size.
@@ 2.2 Minimum moment aberration
For a large number of factors, comparing designs in terms of minimum aberration can become computationally expensive. This is because the computational complexity of the WLP is $O(n2^{m})$, where $n$ is the design’s run size. To overcome this issue, xu2003minimum introduced minimum moment aberration to compare two-level designs. This criterion is equivalent to minimum aberration but uses the moment aberration pattern, a vector based on the similarity between the design’s runs. Let $\mathbf{d}_{i}$ and $\mathbf{d}_{j}$ be the $i$-th and $j$-th run of a $2^{m-p}$ design. We consider $\delta(\mathbf{d}_{i},\mathbf{d}_{j})$ as a function that counts the number of entries in which $\mathbf{d}_{i}$ and $\mathbf{d}_{j}$ coincide. The $t$-th moment of the $2^{m-p}$ design is
@@
In statistical terms, $K_{t}$ is the $t$-th non-central moment of the empirical distribution of similarities between two runs of the design.
The moment aberration pattern is $(K_{1},K_{2},\ldots,K_{m})$, which has a computational complexity of $O(n^{2}m^{2})$ (xu2003minimum). This complexity is much less than that of the WLP when $m$ is large. For two-level designs, xu2003minimum shows a lower bound $K^{\prime}_{i}$ for each element of the moment aberration pattern, which can be used to compute the resolution. That is, the resolution of a design is the smallest value of $i$ for which $K_{i}>K^{\prime}_{i}$.
Similar to minimum aberration, the minimum moment aberration criterion is to minimize the moment aberration pattern from left to right. The moment aberration pattern of the $2^{7-3}$ design in Table 1 is (3.27, 11.67, 42.47, 157.27, 591.27, 2251.67, 8666.47). Since this design has minimum aberration, it also has minimum moment aberration among all designs of the same size.
@@ 3 Methodology
Our methodology to construct two-level fractional factorial designs using LLMs has two main elements. The first one is a prompt template with precise instructions to construct designs, which we develop using prompting techniques (white2023prompt; schulhoff2024prompt). The second one is a computational setup for interacting with the LLMs, which involves APIs and dedicated Python packages. We present each of these elements separately.
@@ 3.1 Prompting techniques
The input of an LLM is a prompt, the writing of which has a critical influence on the quality of the generated output (schulhoff2024prompt). To this end, there are several prompting techniques for crafting prompts, such as role, context, chain of thought, output format, and examples or shots (kojima2022large; white2023prompt; schulhoff2024prompt; sahoo2024systematic). We use these techniques to develop our prompt template, except for the last one for three reasons. First, our template attempts to test the general knowledge of LLMs for constructing two-level fractional factorial designs from scratch using a single well-crafted instruction. Second, writing good examples of two-level fractional factorial designs and their construction requires profound DoE knowledge and expertise, which we do not assume our practitioner has. Third, adding examples to our prompt will significantly increase its length, thereby making it more expensive to run in our APIs; see Section 3.2. We comment on the benefits and challenges of prompts with examples, called few-shot prompts in the prompting literature, in Section 5. Since our prompt template does not include examples, it belongs to the class of zero-shot prompts, which are relevant in practice (schulhoff2024prompt; kojima2022large).
@@ 3.1.1 Role
The performance of an LLM for a given task can be improved by assigning its role in the prompt (wang2024rolellm). This allows the LLM to focus on its expertise in the given role to provide the output. The technique of assigning a role (or persona) to an LLM in the prompt is called role prompting (wang2024rolellm). arvidsson2023prompt, kong2023better and hadi2023large discuss applications of role prompting in engineering, mathematics, and legal practice, respectively.
To construct two-level fractional factorial designs, we assign the role of expert in DoE to the LLM. Specifically, we start our prompt template with the sentence: “You are an expert in the subfield of statistics called design of experiments.” In this role, we make explicit that DoE is a subfield of statistics, so that the LLM concentrates on the experimental design techniques discussed here. In other words, we expect the LLM to focus on two-level fractional factorial designs, resolution, and minimum aberration, which are standard topics in statistical DoE (smucker2022; vazquez2025review).
@@ 3.1.2 Context
The prompt given to an LLM must have the context needed to solve a given problem. This allows the LLM to generate a high-quality solution to the problem. If the prompt lacks context, the LLM might underachieve in the problem or generate a less informative output (sahoo2024systematic; schulhoff2024prompt).
@@
Second, we state the desired number of factors, run size, and levels in the context given in our prompt template. Specifically, we add the following sentences to it: “The number of factors is $m$ and the number of runs is $n$. The factors will be studied at two levels, which are coded as ‘-1’ and ‘1’.” These sentences state that the coded levels of the factors must be 1-1 and 1, which allow us to evaluate the resulting design using our criteria in Section 2. In the sentences, the symbols $m$ and $n$ must be replaced with the desired number of factors and runs, respectively.

### v2

@@ Abstract
Two-level fractional factorial designs permit the study of multiple factors using a limited number of runs. Traditionally, these designs are obtained from catalogs available in standard textbooks or statistical software. However, modern Large Language Models (LLMs) can now produce two-level fractional factorial designs, but the quality of these designs has not been previously assessed. In this paper, we perform a systematic evaluation of two popular classes of LLMs, namely GPT and Gemini models, to construct two-level fractional factorial designs with 8, 16, and 32 runs, and 4 to 26 factors. To this end, we use prompting techniques to develop a high-quality set of design construction tasks for the LLMs. We compare the designs obtained by the LLMs with the best-known designs in terms of resolution and minimum aberration criteria. We show that the LLMs can effectively construct optimal 8-, 16-, and 32-run designs with up to eight factors.
@@ 1 Introduction
Two-level fractional factorial designs (Wu and Hamada, 2011, ch 5) are commonly used experimental plans that can study many factors using an economical number of runs. In these designs, all factors are set at two levels, allowing the study of their main effects and interactions. We denote a two-level fractional factorial design as $2^{m-p}$, where $m$ is the number of factors and $p$ is the number of generated factors. The generated factors are obtained from the interactions of b=mpb=m-p basic factors, whose $2^{b}$ level combinations are all in the design (Wu and Hamada, 2011). Design $2^{m-p}$ is thus a $2^{p}$ fraction of the $2^{m}$ full factorial design, with a run size that is a power of two.
@@
The DoE literature has complete catalogs of two-level fractional factorial designs. For instance, Chen et al. (1993) provide a catalog of designs with 8, 16, 32, and 64 runs, and up to 32 factors. Ryan and Bulutoglu (2010) provide a catalog of 128-run designs with up to 64 factors. The best designs in these catalogs are available in standard DoE textbooks such as Wu and Hamada (2011) and Montgomery (2017), or statistical software such as JMP, Minitab, and the FrF2 package (Grömping, 2014) in R.
@@
In recent years, generative artificial intelligence (GenAI) chatbots, such as ChatGPT (OpenAI, 2025a) and Gemini (Google, 2025b), have emerged as disruptive technology that has transformed the way we learn, write, access information, make decisions, and even conduct research. For example, GenAI chatbots have been used to solve optimization problems (Bertsimas and Margaritis, 2024), prove mathematical theorems (Frieder et al., 2023), teach statistics (Ellis and Slade, 2023), generate code for data analyses (Song et al., 2025), and build search engine software for statistical quality control (Megahed et al., 2024). The success of GenAI chatbots is due to their use of Large Language Models (LLMs; Alammar and Grootendorst, 2024) and the high computational power available these days (TOP500 Project, 2025). Technically, an LLM is a deep neural network (Goodfellow et al., 2016) with billions of parameters that are estimated through a training process that uses an enormous amount of text data. These data are typically extracted from the Internet and may include books, blogs, research and news articles, social media, among others. The goal of the training process is to predict the next word in a sentence. To this end, the LLM uses a neural network architecture called transformer (Vaswani et al., 2017), which allows it to learn the context and relationships between words in the text during its training.
@@
In particular, ChatGPT and Gemini use the Generative Pre-trained Transformer (GPT; OpenAI, 2025a) model and the Gemini model (Comanici et al., 2025), respectively, which are sophisticated, pre-trained transformer-based LLMs. To perform a specific task, their main input (as well as of other LLMs) is a user-specified text called prompt (Schulhoff et al., 2024), which describes the question, instruction, or guidelines for it. At the time of writing this article, the versions of the LLMs used by default in ChatGPT’s and Gemini’s web interfaces are GPT-5.1 and Gemini 2.5 Flash, respectively. Remarkably, these LLMs can reason across audio, vision, and text in real time. Moreover, they can be accessed using application programming interfaces (APIs) for experimentation.
@@
ChatGPT’s output 1 shows a $2^{4-1}$ design that matches the 4-factor 8-run design in Table 8.14 of Montgomery (2017). The four factors are labeled $A$, $B$, $C$, and $D$, where the first three are basic factors and the last is the generated factor obtained as $D=ABC$. That is, the column for $D$ is the element-wise product of the columns that accommodate $A$, $B$, and $C$, when their levels are coded as 1-1 and $1$. The design matrix is shown in the Design Table section. Its first column shows labels for the runs and the other columns show the factors, whose levels are labeled using the symbols ‘-’ and ‘++’. This $2^{4-1}$ design can estimate the main effects and two-factor interactions of the four factors simultaneously. In fact, it is the best design in terms of the resolution and minimum aberration criteria introduced in Section 2. The Appendix shows a similar interaction between us and Gemini with the Gemini 2.5 Flash model, in which the same $2^{4-1}$ design was produced.
Our interactions with ChatGPT and Gemini show that their default LLMs have the potential to construct two-level fractional factorial designs. Therefore, practitioners may feel tempted to construct these designs from scratch using these chatbots, bypassing statistical software and textbooks. To explore the capabilities and limitations of this approach, we conducted a systematic study to evaluate the performance of the GPT-5.1 and Gemini 2.5 Flash models on 36 tasks concerning the construction of two-level fractional factorial designs. The tasks involve designs with 8, 16, and 32 runs, and 4 to 26 factors. Each task has a prompt to instruct the LLM to construct a design. To this end, we develop a prompt template using prompting techniques such as role, context, chain of thought, and output format (Kojima et al., 2022; White et al., 2023; Schulhoff et al., 2024). Our prompt template belongs to the class of zero-shot prompts, introduced in Section 3, and is intended to test the general knowledge of LLMs on constructing two-level fractional factorial designs without examples.
We evaluate the computational performance of the LLMs using 10 replicates of each task. We show that the LLMs have the potential to construct optimal designs with up to 15 factors in terms of the criteria in Section 2. However, the Gemini 2.5 Flash model is better than the GPT-5.1 model because it has a high consistency in obtaining optimal 8-run designs, optimal 16-run designs with up to eight factors, and the optimal 32-run 6-factor design. For designs of other sizes, the performance of both LLMs deteriorates because they cannot consistently construct the optimal design or cannot construct a design at all.
@@ 2 Criteria for two-level fractional factorial designs
We review resolution, minimum aberration, and minimum moment aberration (Wu and Hamada, 2011; Xu, 2003) to evaluate two-level fractional factorial designs. We illustrate these criteria using a $2^{7-3}$ design. Its basic factors are $A$, $B$, $C$, and $D$, and the generated factors are obtained as $E=ABC$, $F=ABD$, and $G=ACD$. The defining relation of the design is
@@ 2.1 Resolution and minimum aberration
The resolution is the length of the shortest word in the defining relation (Wu and Hamada, 2011). It provides a general summary of the properties of a $2^{m-p}$ design for studying main effects and two-factor interactions, which are the most important according to the effect hierarchy principle (Wu and Hamada, 2011, ch 4). For instance, resolution-3 designs provide main effects that are not aliased with each other, but at least one main effect is aliased with one or more two-factor interactions. In contrast, resolution-4 designs provide main effects that are not aliased with each other nor with two-factor interactions. However, at least one pair of two-factor interactions is aliased in these designs. Designs with a resolution equal to five (or higher) are optimal for studying main effects and two-factor interactions because they avoid aliasing among these effects. So, when comparing designs in terms of their resolutions, the design with the higher resolution is preferred. From Equation (1), we have that the $2^{7-3}$ design has a resolution of four.
The minimum aberration criterion allows us to differentiate between designs with the same resolution. It is based on the word length pattern (WLP; Wu and Hamada, 2011), which is a vector that collects the number of words of length 1 to $m$ in the defining relation. Specifically, this vector is $(A_{1},A_{2},\ldots,A_{m})$ where $A_{i}$ is the number of words of length $i$ in the defining relation. Note that the resolution of a design is the smallest value of $i$ for which $A_{i}>0$. From Equation (1), we have that the WLP of the $2^{7-3}$ design is $(0,0,0,7,0,0,0)$. So, the defining relation of this design has seven words of length four.
The minimum aberration criterion is to sequentially minimize the WLP from left to right. In this way, the criterion favors designs with less aliasing involving low-order effects of the factors than high-order ones. Chen et al. (1993) show that the $2^{7-3}$ design in Table 1 has minimum aberration among all two-level fractional factorial designs of the same size.
@@ 2.2 Minimum moment aberration
For a large number of factors, comparing designs in terms of minimum aberration can become computationally expensive. This is because the computational complexity of the WLP is $O(n2^{m})$, where $n$ is the design’s run size. To overcome this issue, Xu (2003) introduced minimum moment aberration to compare two-level designs. This criterion is equivalent to minimum aberration, but it uses the moment aberration pattern, which is a vector based on the similarity between the design’s runs.
Let $\mathbf{d}_{i}$ and $\mathbf{d}_{j}$ be the $i$-th and $j$-th run of a design. We consider $\delta(\mathbf{d}_{i},\mathbf{d}_{j})$ as a function that counts the number of entries in which $\mathbf{d}_{i}$ and $\mathbf{d}_{j}$ coincide. The $t$-th power moment of a two-level design is
@@
In statistical terms, $K_{t}$ is the $t$-th non-central moment of the empirical distribution of similarities between two runs of the design. The moment aberration pattern is $(K_{1},K_{2},\ldots,K_{m})$, which has a computational complexity of $O(n^{2}m^{2})$ (Xu, 2003). This complexity is much lower than that of the WLP when $m$ is large.
Similarly to minimum aberration, the minimum moment aberration criterion is to minimize the moment aberration pattern from left to right. The moment aberration pattern of the $2^{7-3}$ design in Table 1 is (3.27, 11.67, 42.47, 157.27, 591.27, 2251.67, 8666.47). Since this design has minimum aberration, it also has minimum moment aberration among all comparable designs.
In addition to its computational complexity, another advantage of minimum moment aberration is that it can evaluate two-level designs other than $2^{m-p}$ designs. For example, we can use it to evaluate two-level nonregular designs (Wu and Hamada, 2011, ch 8), which are not constructed using basic and generated factors and so, do not have a defining relation. For this reason, we use minimum moment aberration in our numerical evaluations in Section 4 to handle all possible types of two-level designs that LLMs may construct.
In our evaluations, we also use the moment aberration pattern to compute the resolution of a two-level design. To this end, we use the lower bounds for the power moments of a two-level design of Xu (2003). Let $s$ be the smallest value of $i$ for which the $i$-th power moment of the design is equal to its lower bound $K^{\prime}_{i}$. Xu (2003) shows that the design projects into an equally replicated full $2^{s}$ factorial design in every subset of $s$ factors. In other words, the design has a strength of $s$ (Hedayat et al., 2012). Note that, if the first power moment is higher than $K^{\prime}_{1}$, $s=0$ and the design is not level-balanced. Following mee2017selecting, we calculate the resolution of a two-level design as $s+1$. This calculation works for any two-level design and, for $2^{m-p}$ designs, it gives the length of the shortest word in the defining relation.
@@ 3 Methodology
Our methodology to construct two-level fractional factorial designs using LLMs has two main elements. The first one is a prompt template with precise instructions to construct designs, which we develop using prompting techniques (White et al., 2023; Schulhoff et al., 2024). The second one is a computational setup for interacting with the LLMs, which involves APIs and dedicated Python packages. We present each of these elements separately.
@@ 3.1 Prompting techniques
The input of an LLM is a prompt, the writing of which has a critical influence on the quality of the generated output (Schulhoff et al., 2024). To this end, there are several prompting techniques for crafting prompts, such as role, context, chain of thought, output format, and examples or shots (Kojima et al., 2022; White et al., 2023; Schulhoff et al., 2024; Sahoo et al., 2024). We use these techniques to develop our prompt template, except for the last one for three reasons. First, our template attempts to test the general knowledge of LLMs for constructing two-level fractional factorial designs from scratch using a single well-crafted instruction. Second, writing good examples of two-level fractional factorial designs and their construction requires profound DoE knowledge and expertise, which we do not assume our practitioner has. Third, adding examples to our prompt will significantly increase its length, making it more expensive to run in our APIs; see Section 3.2. We comment on the benefits and challenges of prompts with examples, called few-shot prompts in the prompting literature, in Section 5. Since our prompt template does not include examples, it belongs to the class of zero-shot prompts, which are relevant in practice (Schulhoff et al., 2024; Kojima et al., 2022).
@@ 3.1.1 Role
The performance of an LLM for a given task can be improved by assigning its role in the prompt (Wang et al., 2024). This allows the LLM to focus on its expertise in the given role to provide the output. The technique of assigning a role (or persona) to an LLM in the prompt is called role prompting (Wang et al., 2024). Ronanki et al. (2025), Kong et al. (2023) and Hadi et al. (2023) discuss applications of role prompting in engineering, mathematics, and legal practice, respectively.
To construct two-level fractional factorial designs, we assign the role of expert in DoE to the LLM. Specifically, we start our prompt template with the sentence: “You are an expert in the subfield of statistics called design of experiments.” In this role, we make explicit that DoE is a subfield of statistics, so that the LLM concentrates on the experimental design techniques discussed here. In other words, we expect the LLM to focus on two-level fractional factorial designs, resolution, and minimum aberration, which are standard topics in statistical DoE (Smucker et al., 2023; Vazquez and Xuan, 2025).
@@ 3.1.2 Context
The prompt given to an LLM must have the context needed to solve a given problem. This allows the LLM to generate a high-quality solution to the problem. If the prompt lacks context, the LLM might underachieve in the problem or generate a less informative output (Sahoo et al., 2024; Schulhoff et al., 2024).
@@
Second, we state the desired number of factors, run size, and levels in the context given in our prompt template. Specifically, we add the following sentences to it: “The number of factors is $m$ and the number of runs is $n$. The factors will be studied at two levels, which are coded as ‘-1’ and ‘1’.” These sentences state that the coded levels of the factors must be 1-1 and 1, which is standard in DoE. In the sentences, the symbols $m$ and $n$ must be replaced with the desired number of factors and runs, respectively.